# TIT/HEP-708

### v1

@@
TBA and wall-crossing for deformed supersymmetric quantum mechanics
@@
We study the deformed supersymmetric quantum mechanics with a polynomial superpotential with $\hbar$-correction. In the minimal chamber, where all turning points are real and distinct, it was shown that the exact WKB periods obey the ${\mathbb{Z}}_{4}$-extended thermodynamic Bethe ansatz (TBA) equations of the undeformed potential. By changing the energy parameter above/below the critical points, the turning points become complex, and the moduli are outside of the minimal chamber. We study the wall-crossing of the ${\mathbb{Z}}_{4}$-extended TBA equations by this change of moduli and show that the ${\mathbb{Z}}_{4}$-structure is preserved after the wall-crossing. In particular, the TBA equations for the cubic superpotential are studied in detail, where there are two chambers (minimal and maximal). At the maximally symmetric point in the maximal chamber, the TBA system becomes the two sets of the $D_{3}$-type TBA equations, which are regarded as the ${\mathbb{Z}}_{4}$-extension of the $A_{3}/{\mathbb{Z}}_{2}$-TBA equation.
@@ 1 Introduction
The connection between exact WKB periods and TBA equations was initially established in one-dimensional quantum mechanics with a polynomial potential. The corresponding Schrödinger equation can also be interpreted as the quantum Seiberg-Witten (SW) curve of $A_{r}$-type Argyres-Douglas (AD) theory in the Nekrasov–Shatashvili (NS) limit of the Omega background [7, 8]. In this interpretation, the WKB periods correspond to central charges of BPS particles in the AD theory. The corresponding TBA equations can also be obtained by taking the conformal limit of the non-linear integral equations for the central charges of the BPS particles in the Gaiotto–Moore–Neitzke (GMN) formalism [9, 10], which will be called GMN TBA in this paper.
@@
Recently, the resurgence theory, e.g. the exact WKB method, was applied to the deformed supersymmetric quantum mechanics[27, 28, 29]. The model includes a bosonic field and some Grassmann fields with a superpotential. After integrating out the fermionic degrees of freedom, the resulting model has an effective potential with $\hbar$-correction. In the previous work [30], we studied the relation between the exact WKB periods in the deformed supersymmetric quantum mechanics and the TBA equations. A key distinction from earlier studies is that this Schrödinger equation does not appear as a quantum SW curve, which suggests it may lie outside the scope of the conventional GMN formalism. Applying the exact WKB method and the ODE/IM correspondence to this model, we derived a set of $\mathbb{Z}_{4}$-extended TBA equations. A notable feature of the effective potential is the emergence of a $\hbar$ corrected term, which distinguishes it significantly from the standard exact WKB analysis for the Schrödinger equation with only classical potential. In particular, odd power terms will appear in the $\hbar$ expansion of the WKB periods. In this case, the discontinuity structure of the WKB periods can be understood from the $\mathbb{Z}_{4}$-extended TBA equations. For the cubic superpotential with real and distinct turning points, these TBA equations split into two independent $D_{3}$-type TBA equations. Notably, these $D_{3}$-type TBA systems also appear as the GMN TBA equations of the $D_{3}$-type AD theory in the minimal chamber. It is thus interesting to extend the TBA equations to the whole moduli space of the superpotential, where the TBA equations will be modified at the marginal stability wall of the WKB periods (BPS particles). This modification is called the wall-crossing of the TBA equations.
In this paper, we present a detailed study on the wall-crossing of the $\mathbb{Z}_{4}$-extended TBA equations to explore the correspondence between the exact WKB method and the quantum integrable models for the deformed supersymmetric quantum mechanics. This paper is organized as follows: In section 2, we review the exact WKB analysis for the Schrödinger equation with $\hbar$-deformed potential. We then present the $\mathbb{Z}_{4}$-extended TBA equations, and show the general procedure of the wall-crossing of the $\mathbb{Z}_{4}$-extended TBA in section 3. In section 4, we will apply the general procedure to the cubic superpotential. We work out in detail the monomial superpotential, where the effective potential becomes the quartic monomial potential with $\hbar$-correction.
@@ 2 Exact WKB Analysis for the Schrödinger equation
In this section, we briefly review the exact WKB analysis for the Schrödinger equation for the effective potential with $\hbar$-correction:
@@
We now study the WKB solution of Eq. (1), where we use the ansatz:
@@
For the study of the spectral problem, it is important to consider the period integral $\oint P(x)dx$ on the WKB curve $\Sigma_{\rm WKB}$:
@@
For the superpotential $W(x)$ in (3), $Q_{0}(x)$ is a polynomial of order $2N$. Then it has $2N$ zeros, denoted as $x_{1},\cdots,x_{2N}$, which are called turning points. First, we assume that these turning points are all real, distinct, and ordered $x_{1}<\cdots<x_{2N}$. We denote the one-cycle going around the interval $x_{s},x_{s+1}]$ by $\gamma_{s}$ ($s=1,\cdots,2N-1$). The one-cycle $\gamma_{s}$ with even $s$ (odd $s$) is also assumed to correspond to the forbidden (allowed) region, where the potential term $Q_{0}(x)$ becomes positive (negative). We will choose the branch cut and the orientation of each cycle such that the corresponding classical period, $\oint_{\gamma_{s}}p_{0}(x)dx$, is positive real (positive imaginary) for the forbidden cycle (allowed cycle). For a general WKB curve, periods can be obtained by deforming the locations of the turning points.
@@
$\Pi_{\gamma}^{(n)}=\sum_{k=0}^{2N-2}B_{k}^{(n)}\widehat{\Pi}_{\gamma,k},$ (8)
where $B_{k}^{(n)}$ are constants and $\widehat{\Pi}_{\gamma,k}$ are the period integrals of the basis of meromorphic differentials on the WKB curve:
$\widehat{\Pi}_{\gamma,k}=\oint_{\gamma}\frac{x^{k}}{y}dx,\quad k=0,\dots,2N-2.$ (9)
It is known that the $\hbar$ expansion (7) is an asymptotic series. It is treated as an analytic function after the Borel transform:
@@
The lateral Borel resummation shows discontinuities of the period and characterizes the analytic structure of the exact WKB periods. For a polynomial potential without $\hbar$-correction, the discontinuities are determined by the Delabaere-Dillinger-Pham (DDP) formula [33, 34, 35]. It turns out the DDP formula can be expressed in the form of the non-linear integral equations called the TBA equations, which have been studied in the context of BPS spectra in ${\cal N}=2$ supersymmetric gauge theories [9, 10]. In the case of a polynomial potential in the minimal chamber, which includes the case where all the turning points are real and different, the TBA equations take the form of the $A_{r}$-type. However, the structure of TBA equations changes when potential parameters are outside the minimal chamber. This is called the wall-crossing phenomenon.
For deformed supersymmetric quantum mechanics, the TBA equations in the minimal chamber take the form of ${\mathbb{Z}}_{4}$-extension of the undeformed ones. The main purpose of the present work is to explore the wall-crossing of the ${\mathbb{Z}}_{4}$-extended TBA equations for the deformed supersymmetric quantum mechanics which includes the $\hbar$-correction in the potential.
3 $\mathbb{Z}_{4}$-extended TBA and wall-crossing
In this section, we first present the $\mathbb{Z}_{4}$-extended TBA equations in the minimal chamber for the deformed supersymmetric quantum mechanics with superpotential (3) using the ODE/IM correspondence [30]. We then provide a general procedure for wall-crossing of these $\mathbb{Z}_{4}$-extended TBA to extend them to the whole space of moduli parameters.
@@ 3.1 $\mathbb{Z}_{4}$-extended TBA from the ODE/IM correspondence
We first study the asymptotic solutions of the ODE at infinity and the Y-system constructed from the Wronskians of the solutions. For positive real $\hbar$, the fastest decaying solution of (1) at infinity around the positive real axis is
@@
where $n_{N}=-\frac{N}{2}-B_{N+1}$. Here, $B_{n}$ is defined by
@@
$Y_{s}(\hbar,e^{a\frac{\pi i}{2}}m)=Y_{a,s}(\hbar).$ (20)
We first focus on the minimal chamber. In this case, we can identify the one-cycle $\gamma_{s}$ on the WKB curve with the Y-function $Y_{s}$ [36]. Let us introduce the spectral parameter $\theta=-\log\hbar$. Then, from the analyticity in the strip $|{\rm Im}\theta|<{\pi\over 2}$ and the asymptotic behavior $Y_{s}(\theta)\rightarrow-m_{s}e^{\theta}$ at $\theta\rightarrow+\infty$, the Y-system can be converted into the $\mathbb{Z}_{4}$-extended TBA equations
@@
Note that $m^{(\frac{1}{2})}$ plays the role of chemical potential in the TBA equations. Setting $m=0$, $Y_{a,s}$ are identified for any $a=0,\cdots,3$, which leads to the original $A_{2N-1}$-type TBA equations [6].
Compared to the undeformed TBA in [6], the kernel $K(\theta)=1/(2\pi\cosh(\theta))$ is decomposed into two kernels $K_{\pm}$. Due to the $\tanh(\theta)$ term, the kernels approach to nonzero constants at large $\theta$, which will cause difficulty in the numerical study. Expanding the kernels at large $\theta$, the term $\tanh(\theta)$ in the kernel will also lead to odd power terms in the $\hbar$-expansion of WKB periods, which do not appear in the Schrödinger equation without $\hbar$-deformed potential. Moreover, singularities will appear in the kernel of TBA equations if we shift $\theta$ by $\pm\pi i/2$, which leads to discontinuity of $\log Y(\theta\pm\pi i/2)$ [^4] This will lead to the DDP formula for the corresponding WKB periods. Very different from the original TBA, the DDP formula differs in the direction ${\rm Im}(\theta)=\pi/2$ and ${\rm Im}(\theta)=-\pi/2$..
@@
It is critical to note that poles of the kernel $K_{\pm}(\theta-\theta^{\prime}-i\phi_{s}+i\phi_{s\pm 1})$, i.e. $\theta^{\prime}=\theta-i\phi_{s}+i\phi_{s\pm 1}\pm i\pi/2$, will collide with the real and positive axis of the convolution as $|\phi_{s}-\phi_{s\pm 1}|=\pi/2$, where one has to modify the TBA equations by picking up the contribution of poles. TBA equations (24) are only valid in the regions $|\phi_{s}-\phi_{s\pm 1}|<\pi/2$, which is denoted as the minimal chamber. In the following section, we will provide the general procedure of the wall-crossing for the $\mathbb{Z}_{4}$-extended TBA.
3.2 Wall-crossing of $\mathbb{Z}_{4}$-extended TBA
Without loss of generality, we consider the wall-crossing of TBA equations when $\phi_{2}-\phi_{1}$ increases below $\pi/2$ and goes above. In this case, we should pick up the contribution of the poles of $K_{\pm;1,2}(\theta-\theta^{\prime})$ and $K_{\pm;2,1}(\theta-\theta^{\prime})$, where the poles are located at $\theta^{\prime}=\theta-i\phi_{1}+i\phi_{2}-\frac{i\pi}{2}$ and $\theta^{\prime}=\theta-i\phi_{2}+i\phi_{1}+\frac{i\pi}{2}$, respectively. Picking up these contributions, we obtain the TBA equations
@@ 3.2 Wall-crossing of $\mathbb{Z}_{4}$-extended TBA
The summation of the left-hand leads to
@@
The denominator can be cancelled by the terms of the summation of the right-hand side, which finally leads to
@@
where we have shifted the argument of the integrals associated with $Y_{12}^{\rm n}$. This provides the definition of the effective central charge for TBA equations (34) in the context of quantum integrable models. It is critical to note that the value of the effective central charge is invariant under the process of wall-crossing. The wall-crossing of ${\mathbb{Z}}_{4}$-extended TBA associated with other poles of the kernel can be evaluated similarly.

### v2

@@
Thermodynamic Bethe ansatz and wall crossing for deformed supersymmetric quantum mechanics
@@
We study the deformed supersymmetric quantum mechanics with a polynomial superpotential with $\hbar$ correction. In the minimal chamber, where all turning points are real and distinct, it was shown that the exact Wentzel–Kramers–Brillouin periods obey the ${\mathbb{Z}}_{4}$-extended thermodynamic Bethe ansatz (TBA) equations of the undeformed potential. By changing the energy parameter above/below the critical points, the turning points become complex, and the moduli are outside of the minimal chamber. We study the wall crossing of the ${\mathbb{Z}}_{4}$-extended TBA equations by this change of moduli and show that the ${\mathbb{Z}}_{4}$ structure is preserved after the wall crossing. In particular, the TBA equations for the cubic superpotential are studied in detail, where there are two chambers (minimal and maximal). At the maximally symmetric point in the maximal chamber, the TBA system becomes the two sets of the $D_{3}$-type TBA equations, which are regarded as the ${\mathbb{Z}}_{4}$ extension of the $A_{3}/{\mathbb{Z}}_{2}$-type TBA equation.
@@ 1 Introduction
The connection between exact WKB periods and TBA equations was initially established in one-dimensional quantum mechanics with a polynomial potential. The corresponding Schrödinger equation can also be interpreted as the quantum Seiberg–Witten (SW) curve of $A_{r}$-type Argyres–Douglas (AD) theory in the Nekrasov–Shatashvili (NS) limit of the omega background [7, 8]. In this interpretation, the WKB periods correspond to central charges of BPS particles in the AD theory. The corresponding TBA equations can also be obtained by taking the conformal limit of the non-linear integral equations for the central charges of the BPS particles in the Gaiotto–Moore–Neitzke (GMN) formalism [9, 10], which will be called GMN TBA in this paper.
@@
Recently, the resurgence theory, e.g. the exact WKB method, was applied to the deformed supersymmetric quantum mechanics[27, 28, 29]. The model includes a bosonic field and some Grassmann fields with a superpotential. After integrating out the fermionic degrees of freedom, the resulting model has an effective potential with $\hbar$ correction. In the previous work [30], we studied the relation between the exact WKB periods in the deformed supersymmetric quantum mechanics and the TBA equations. A key distinction from earlier studies is that this Schrödinger equation does not appear as a quantum SW curve, which suggests it may lie outside the scope of the conventional GMN formalism. Applying the exact WKB method and the ODE/IM correspondence to this model, we derived a set of $\mathbb{Z}_{4}$-extended TBA equations. A notable feature of the effective potential is the emergence of an $\hbar$ corrected term, which distinguishes it significantly from the standard exact WKB analysis for the Schrödinger equation with only classical potential. In particular, odd power terms will appear in the $\hbar$ expansion of the WKB periods. In this case, the discontinuity structure of the WKB periods can be understood from the $\mathbb{Z}_{4}$-extended TBA equations. For the cubic superpotential with real and distinct turning points, these TBA equations split into two independent $D_{3}$-type TBA equations. Notably, these $D_{3}$-type TBA systems also appear as the GMN TBA equations of the $D_{3}$-type AD theory in the minimal chamber. It is thus interesting to extend the TBA equations to the whole moduli space of the superpotential, where the TBA equations will be modified at the marginal stability wall of the WKB periods (BPS particles). This modification is called the wall crossing of the TBA equations.
In this paper, we present a detailed study on the wall crossing of the $\mathbb{Z}_{4}$-extended TBA equations to explore the correspondence between the exact WKB method and the quantum integrable models for the deformed supersymmetric quantum mechanics. This paper is organized as follows: In section 2, we review the exact WKB analysis for the Schrödinger equation with $\hbar$-deformed potential. We then present the $\mathbb{Z}_{4}$-extended TBA equations, and show the general procedure of the wall crossing of the $\mathbb{Z}_{4}$-extended TBA in section 3. In section 4, we will apply the general procedure to the cubic superpotential. We work out in detail the monomial superpotential, where the effective potential becomes the quartic monomial potential with $\hbar$ correction.
@@ 2 Exact WKB Analysis for the Schrödinger equation
In this section, we briefly review the exact WKB analysis for the Schrödinger equation for the effective potential with $\hbar$ correction:
@@
Here $m$ and $u_{j}$ can be any complex number in general. We now study the WKB solution of Eq. (1), where we use the ansatz:
@@
For the study of the spectral problem, it is important to consider the period integral $\oint P(x)dx$ on the WKB curve $\Sigma_{\rm WKB}$, which is defined by the algebraic relation of complex coordinates $(y,x)$:
@@
The solution of $y$ provides the leading order term $p_{0}$ of $P(x)$. For the superpotential $W(x)$ in (3), $Q_{0}(x)$ is a polynomial of order $2N$. Then it has $2N$ zeros, denoted as $x_{1},\cdots,x_{2N}$, which are called turning points. First, we assume that these turning points are all real, distinct, and ordered $x_{1}<\cdots<x_{2N}$. We denote the one-cycle going around the interval $x_{s},x_{s+1}]$ by $\gamma_{s}$ ($s=1,\cdots,2N-1$). The one-cycle $\gamma_{s}$ with even $s$ (odd $s$) is also assumed to correspond to the forbidden (allowed) region, where the potential term $Q_{0}(x)$ becomes positive (negative). We will choose the branch cut and the orientation of each cycle such that the corresponding classical period, $\oint_{\gamma_{s}}p_{0}(x)dx$, is positive real (positive imaginary) for the forbidden cycle (allowed cycle). For a general WKB curve, periods can be obtained by deforming the locations of the turning points.
@@
$\Pi_{\gamma}^{(n)}=\sum_{k=0}^{2N-2}B_{k}^{(n)}\big(\oint_{\gamma}\frac{x^{k}}{y}dx\big),\quad k=0,\dots,2N-2,$ (8)
where $B_{k}^{(n)}$ are constants and $\oint_{\gamma}\frac{x^{k}}{y}dx$ are the period integrals of the basis of meromorphic differentials on the WKB curve. It is known that the $\hbar$ expansion (7) is an asymptotic series. It is treated as an analytic function after the Borel transform:
@@
The lateral Borel resummation shows discontinuities of the period and characterizes the analytic structure of the exact WKB periods. For a polynomial potential without $\hbar$ correction, the discontinuities are determined by the Delabaere–Dillinger–Pham (DDP) formula [33, 34, 35]. It turns out the DDP formula can be expressed in the form of the non-linear integral equations called the TBA equations, which have been studied in the context of BPS spectra in ${\cal N}=2$ supersymmetric gauge theories [9, 10]. In the case of a polynomial potential in the minimal chamber, which includes the case where all the turning points are real and different, the TBA equations take the form of the $A_{r}$-type. However, the structure of TBA equations changes when potential parameters are outside the minimal chamber. This is called the wall crossing phenomenon.
For deformed supersymmetric quantum mechanics, the TBA equations in the minimal chamber take the form of ${\mathbb{Z}}_{4}$ extension of the undeformed ones. The main purpose of the present work is to explore the wall crossing of the ${\mathbb{Z}}_{4}$-extended TBA equations for the deformed supersymmetric quantum mechanics which includes the $\hbar$ correction in the potential.
3 $\mathbb{Z}_{4}$-extended TBA and wall crossing
In this section, we first present the $\mathbb{Z}_{4}$-extended TBA equations in the minimal chamber for the deformed supersymmetric quantum mechanics with superpotential (3) using the ODE/IM correspondence [30]. We then provide a general procedure for wall crossing of these $\mathbb{Z}_{4}$-extended TBA to extend them to the whole space of moduli parameters.
@@ 3.1 $\mathbb{Z}_{4}$-extended TBA from the ODE/IM correspondence
We first study the asymptotic solutions of the ODE at infinity and the Y-system constructed from the Wronskians of the solutions. At infinity, there are two kinds of asymptotic solutions: growing and decaying. Since only the decaying solution is uniquely defined, we will focus on the fastest decaying solution. For positive real $\hbar$, the fastest decaying solution of (1) at infinity $y(x,u_{j},E,m,\hbar)$ around the positive real axis is
@@
where $n_{N}=-\frac{N}{2}-B_{N+1}$. Note that $E$ and $m$ will appear in the subleading order of $y(x,u_{j},E,m,\hbar)$, which can be obtained by solving (1). Here, $B_{n}$ is defined by
@@
$Y_{a,s}(\hbar):=Y_{s}(\hbar,e^{a\frac{\pi i}{2}}m),$ (19)
where we have omitted the argument $m$ on the left hand side. We aim to convert the Y-system into integral equations that depend on the parameters $u_{j},E$, and $m$. We first focus on the minimal chamber. In this case, we can identify the one-cycle $\gamma_{s}$ on the WKB curve with the Y-function $Y_{s}$ [36]. Let us introduce the spectral parameter $\theta=-\log\hbar$. Then, from the analyticity in the strip $|{\rm Im}\theta|<{\pi\over 2}$ and the asymptotic behavior $Y_{s}(\theta)\rightarrow-m_{s}e^{\theta}$ at $\theta\rightarrow+\infty$, the Y-system can be converted into the $\mathbb{Z}_{4}$-extended TBA equations
@@
Note that $m_{a,s}^{(\frac{1}{2})}$ plays the role of chemical potential in the TBA equations. Setting $m=0$, $Y_{a,s}$ for any $a$ are identified, which leads to the $A_{2N-1}$-type TBA equations [6]. Compared to the undeformed TBA in [6], the kernel $K(\theta)=1/(2\pi\cosh(\theta))$ is decomposed into two kernels $K_{\pm}$. Due to the $\tanh(\theta)$ term, the kernels approach to nonzero constants at large $\theta$, which will cause difficulty in the numerical study. Expanding the kernels at large $\theta$, the term $\tanh(\theta)$ in the kernel will also lead to odd power terms in the $\hbar$-expansion of WKB periods, which do not appear in the Schrödinger equation without $\hbar$-deformed potential. Moreover, singularities will appear in the kernel of TBA equations if we shift $\theta$ by $\pm\pi i/2$, which leads to discontinuity of $\log Y(\theta\pm\pi i/2)$ [^4] This will lead to the DDP formula for the corresponding WKB periods. Very different from the original TBA, the DDP formula differs in the directions ${\rm Im}(\theta)=\pi/2$ and ${\rm Im}(\theta)=-\pi/2$..
@@
It is critical to note that poles of the kernel $K_{\pm}(\theta-\theta^{\prime}-i\phi_{s}+i\phi_{s\pm 1})$, i.e. $\theta^{\prime}=\theta-i\phi_{s}+i\phi_{s\pm 1}\pm i\pi/2$, will collide with the real and positive axis of the convolution as $|\phi_{s}-\phi_{s\pm 1}|=\pi/2$, where one has to modify the TBA equations by picking up the contribution of poles. TBA equations (23) are only valid in the regions $|\phi_{s}-\phi_{s\pm 1}|<\pi/2$, which is denoted as the minimal chamber. In the following section, we will provide the general procedure of the wall crossing for the $\mathbb{Z}_{4}$-extended TBA.
3.2 wall crossing of $\mathbb{Z}_{4}$-extended TBA
Without loss of generality, we consider the wall crossing of TBA equations when $\phi_{2}-\phi_{1}$ increases below $\pi/2$ and goes above. In this case, we should pick up the contribution of the poles of $K_{\pm;1,2}(\theta-\theta^{\prime})$ and $K_{\pm;2,1}(\theta-\theta^{\prime})$, where the poles are located at $\theta^{\prime}=\theta-i\phi_{1}+i\phi_{2}-\frac{i\pi}{2}$ and $\theta^{\prime}=\theta-i\phi_{2}+i\phi_{1}+\frac{i\pi}{2}$, respectively. Picking up these contributions, we obtain the TBA equations
@@
The summation of the left-hand side leads to
@@
The denominator can be canceled by the terms of the summation of the right-hand side, which finally leads to
@@
where we have shifted the argument of the integrals associated with $Y_{12}^{\rm n}$. This provides the definition of the effective central charge for TBA equations (33) in the context of quantum integrable models. It is critical to note that the value of the effective central charge is invariant under the process of wall crossing. The wall crossing of ${\mathbb{Z}}_{4}$-extended TBA associated with other poles of the kernel can be evaluated similarly.